# Delving into ChatGPT usage in academic writing through excess vocabulary

### v4

@@ 4.1 Dataset
The PubMed data until the end of 2023 were collected and shared by González-Márquez et al. (2024). They used the annual PubMed snapshot from the end of 2023, and parsed the XML files to extract the abstract texts and some metadata, keeping only complete English-language abstracts with length 250–4000 characters. Their resulting dataset contains 23 389 083 abstracts.
We added to these data all daily PubMed (https://pubmed.ncbi.nlm.nih.gov/download) update files until the end of June 2024, and parsed the XML files in the same way, applying the same filtering criteria as González-Márquez et al. (2024). We kept the last version of the abstract for each unique PMID (PubMed ID), because the daily update files often contain updated information about the papers that were already listed in previous files. This resulted in 24 105 518 abstracts.
We then only analyzed papers with publication years from 2010 to 2024, giving us 14 448 711 abstracts for analysis.
@@ 4.2 Pre-processing
We spent substantial effort to clean the abstracts from all such contaminating strings, using over 100 regular expressions to find and eliminate them. Overall, 270 189 abstracts were affected by our cleaning procedure. We have also entirely erased 3431 abstracts of errata, corrigenda, correction, or retraction notices (identified based on titles).
We then computed a binary word occurrence matrix using CountVectorizer(binary=True) from Scikit-learn, obtaining a sparse matrix. We focused the subsequent analysis on words consisting of at least four letters and composed only out of the 26 letters of English alphabet. Note that different strings (e.g. string and strings) were treated as two distinct words.
@@ 4.3 Statistical analysis
4.4 Word annotations
We identified 829 unique excess words (surpassing thresholds on or ) from 2013 to 2024. Some of these words showed excess usage in multiple years. We sorted the list alphabetically and annotated them as content and style words while being blinded to the year in which they were selected as excess words. We assigned parts of speech (nouns, adjectives, verbs, etc.) in the same way. In case of doubt, the words were discussed between the authors. When we were not certain whether a word was content or style because of ambiguous usage, we did not label this word as either content or style.
@@ 4.5 Subgroup analysis
For the analysis presented in Figure 5 we separately analysed the following subgroups: 50 countries with the most papers in our dataset; 100 journals with the most papers in our dataset in 2024; all 39 domains taken from González-Márquez et al. (2024) (where domains were assigned based on the journal names, e.g., assigning all papers from The Journal of Neuroscience to the ‘neuroscience’ domain); male and female inferred genders of the first and of the last authors (inferred via the gender package, Blevins and Mullen, 2015).
@@
Our gender inference aims to capture perceived gender based on first name and is only approximate. The inference model has clear shortcomings, including limited US-based training data. Moreover, some first names are inherently gender-ambiguous. We were able to infer genders only for 55% of the authors.
@@ 4.6 LLM usage
We did not use ChatGPT or any other LLMs for writing the manuscript or for performing the data analysis.
@@ 4.7 Data and code availability
Our analysis code in Python is available at https://github.com/berenslab/chatgpt-excess-words. The original PubMed data are openly available for bulk download here: https://pubmed.ncbi.nlm.nih.gov/download.
@@ Acknowledgements
This work was conceived at the Dagstuhl seminar 24122 supported by the Leibniz Center for Informatics. This work was funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via Germany’s Excellence Strategy (Excellence cluster 2064 “Machine Learning — New Perspectives for Science”, EXC 390727645), the German Ministry of Science and Education (BMBF) via the Tübingen AI Center (01IS18039A) and the Gemeinnützige Hertie-Stiftung. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting RGM. EÁH’s work is supported by NSF CAREER Grant No IIS-1943506.

### v5

@@ 4.1 Dataset
We used the annual PubMed snapshot released in the beginning of 2025 (https://ftp.ncbi.nlm.nih.gov/pubmed/baseline/), containing files from pubmed25n0001.xml.gz to pubmed25n1274.xml.gz. We parsed the XML files as described in our prior work (González-Márquez et al., 2024), applying the same filtering criteria, to extract the abstract texts and some metadata, keeping only complete English-language abstracts with length 250–4000 characters. This resulted in 24 814 136 abstracts. We then only analyzed papers with publication years from 2010 to 2024, giving us 15 103 888 abstracts for analysis.
For each paper, we defined the affiliation country as the country of the first affiliation of the first author. We assigned papers to the 39 fields taken from González-Márquez et al. (2024) based on the journal names: for example, all papers from all journals containing the word neuroscience (such as The Journal of Neuroscience or Nature Neuroscience) were assigned to the ‘neuroscience’ field. We used the first names of the first and the last authors to infer their genders via the gender package (Blevins and Mullen, 2015). Our gender inference aims to capture perceived gender based on first name and is only approximate. The inference model has clear shortcomings, including limited US-based training data. Moreover, some first names are inherently gender-ambiguous. We were able to infer genders only for 55% of the authors. See González-Márquez et al. (2024) for further details.
@@ 4.2 Pre-processing
We spent substantial effort to clean the abstracts from all such contaminating strings, using over 100 regular expressions to find and eliminate them. Overall, 286 744 abstracts were affected by our cleaning procedure. We have also entirely erased 3 514 abstracts of errata, corrigenda, correction, or retraction notices (identified based on titles).
We then computed a binary word occurrence matrix using CountVectorizer(binary=True, min_df=1e-6) from Scikit-learn (Pedregosa et al., 2011), obtaining a sparse matrix. We focused the subsequent analysis on words consisting of at least four letters and composed only out of the 26 letters of English alphabet. Note that different strings (e.g. mask and masks) were treated as two distinct words.
@@ 4.3 Statistical analysis
4.4 Word annotations and lemmatisation
We identified 900 unique excess words (surpassing our thresholds on or ) from 2013 to 2024. Some of these words showed excess usage in multiple years. We sorted the list alphabetically and annotated them as content and style words while being blinded to the year in which they were selected as excess words. We assigned parts of speech (nouns, adjectives, verbs, etc.) in the same way. In case of doubt, the words were discussed between the authors. When we were not certain whether a word was content or style because of ambiguous usage, we did not label this word as either content or style.
To lemmatise the words for Figure S5, we used WordNetLemmatizer() from the Python NLTK library (Bird et al., 2009), and manually added several lemmas such as chatbotschatbot and circrnascircrna. We also converted British spelling to the US spelling using a dictionary from https://github.com/hyperreality/American-British-English-Translator.
@@ 4.5 Subgroup analysis
For the analysis presented in Figure 5 we separately analysed the following subgroups: 50 affiliation countries with the most papers in our dataset; 100 journals with the most papers in our dataset in 2024; all 39 fields taken from González-Márquez et al. (2024); male and female inferred genders of the first and of the last authors.
@@ 4.6 LLM usage
We did not use any LLMs for writing or editing the manuscript.
@@ 4.7 Data and code availability
Our analysis code in Python is available at https://github.com/berenslab/llm-excess-vocab. The matrix of yearly word occurrences (for each word and year, the number of abstracts in that year containing that word; the additional last row contains the total number of abstracts in that year) is available in our Github repository as a csv.gz file. Our content/style annotations are also available there as a csv file. The original PubMed data are openly available for bulk download here: https://pubmed.ncbi.nlm.nih.gov/download.
Author contributions
Conceived and designed the experiments: all authors. Performed the experiments: DK. Analyzed the data: DK. Contributed materials/analysis tools: DK, RGM. Wrote the paper: all authors.
@@ Acknowledgements
This work was conceived at the Dagstuhl seminar 24122 supported by the Leibniz Center for Informatics. This work was funded by the Deutsche Forschungsgemeinschaft (KO6282/2-1) and the Gemeinnützige Hertie-Stiftung. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting RGM. DK is a member of the Germany’s Excellence cluster 2064 “Machine Learning — New Perspectives for Science” (EXC 390727645). EÁH’s work is supported by NSF CAREER Grant No IIS-1943506. We thank The Economist coverage for inspiring us to use stacked bars in Figure 3.
@@
accentuates, acknowledges, acknowledging, addresses, adept, adhered, adhering, advancement, advancements, advancing, advocates, advocating, affirming, afflicted, aiding, akin, align, aligning, aligns, alongside, amidst, assessments, attains, attributed, augmenting, avenue, avenues, bolster, bolstered, bolstering, broader, burgeoning, capabilities, capitalizing, categorized, categorizes, categorizing, combating, commendable, compelling, complicates, complicating, comprehending, comprising, consequently, consolidates, contributing, conversely, correlating, crafted, crafting, culminating, customizing, delineates, delve, delved, delves, delving, demonstrating, dependability, dependable, detailing, detrimentally, diminishes, diminishing, discern, discerned, discernible, discerning, displaying, disrupts, distinctions, distinctive, elevate, elevates, elevating, elucidate, elucidates, elucidating, embracing, emerges, emphasises, emphasising, emphasize, emphasizes, emphasizing, employing, employs, empowers, emulating, emulation, enabling, encapsulates, encompass, encompassed, encompasses, encompassing, endeavors, endeavours, enduring, enhancements, enhances, ensuring, equipping, escalating, evaluates, evolving, exacerbating, examines, exceeding, excels, exceptional, exceptionally, exerting, exhibiting, exhibits, expedite, expediting, exploration, explores, facilitated, facilitates, facilitating, featuring, formidable, fostering, fosters, foundational, furnish, garnered, garnering, gauged, grappling, groundbreaking, groundwork, harness, harnesses, harnessing, heighten, heightened, hinder, hinges, hinting, hold, holds, illuminates, illuminating, imbalances, impacting, impede, impeding, imperative, impressive, inadequately, incorporates, incorporating, influencing, inherent, initially, innovative, inquiries, integrates, integrating, integration, interconnectedness, interplay, intricacies, intricate, intricately, introduces, invaluable, investigates, involves, juxtaposed, leverages, leveraging, maintaining, merges, methodologies, meticulous, meticulously, multifaceted, necessitate, necessitates, necessitating, necessity, notable, noteworthy, nuanced, nuances, offering, optimizing, orchestrating, outlines, overlook, overlooking, paving, persist, pinpoint, pinpointed, pinpointing, pioneering, pioneers, pivotal, poised, pose, posed, poses, posing, predominantly, preserving, pressing, promise, pronounced, propelling, realm, realms, recognizing, refine, refines, refining, remarkable, renowned, revealing, reveals, revolutionize, revolutionizing, revolves, scrutinize, scrutinized, scrutinizing, seamless, seamlessly, seeks, serves, serving, shaping, shedding, showcased, showcases, showcasing, signifying, solidify, spanned, spanning, spurred, stands, stemming, strategically, streamline, streamlined, streamlines, streamlining, struggle, substantiated, substantiates, surged, surmount, surpass, surpassed, surpasses, surpassing, swift, swiftly, thorough, transformative, typically, ultimately, uncharted, uncovering, underexplored, underscore, underscored, underscores, underscoring, unexplored, unlocking, unparalleled, unraveling, unveil, unveiled, unveiling, unveils, uphold, upholding, urging, utilizes, varying, versatility, warranting, yielding